write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form ax^2+bx+c=0.
Solution: 10, only one solution

Answer by user_dude2008(1862)

(x-10)(x-10)=x^2-10x-10x+100=x^2-20x+100

Answer: polynomial is y=x^2-20x+100